Solve each equation by first eliminating the fractions.
1/2q - 6 = 1/5q

Solve each equation by first eliminating the decimal numbers
0.5b + 3.4 = 0.2b + 12.4

Solve each equation. If you feel proficient enough, try simplying the process
4x - 2 = -8 + 5x

Solve each equation. Identify each as a conditional equation, an inconsistent equation, or a identity
3 - 3(5 - x) = 0

Solve each equation
-3 - 4(t - 5) = -2(t + 3) + 11

Answer by stanbon(75887)

Solve each equation by first eliminating the fractions.
1/2q - 6 = 1/5q
Multiply through by 10q to get:
5-60q=2
60q=3
q=1/20
Solve each equation by first eliminating the decimal numbers
0.5b + 3.4 = 0.2b + 12.4
Multiply throu by 10 to get:
5b+34=2b+124
3b=90
b=20
Solve each equation. If you feel proficient enough, try simplying the process
4x - 2 = -8 + 5x
x=6

Solve each equation. Identify each as a conditional equation, an inconsistent equation, or a identity
3 - 3(5 - x) = 0
3-15+3x=0
3x=12
x=4
Solve each equation
-3 - 4(t - 5) = -2(t + 3) + 11
-3-4t+20=-2t-6+11
2t=-6+11+3-20
2t=-12
t=-6